# Decision Theoretic Foundations for Experiments Evaluating Human Decisions

### v3

@@ 5. Recommendations for Practice
The lack of consensus around how to define a decision problem and what the minimum required components for normative analysis should be leads to research landscape in which researchers can easily be tempted to overinterpret the results of their studies, and readers may not recognize the overinterpretations. Only when researchers define the task they study as a complete decision problem can the results of experiments that study slightly different problems (e.g., use different data-generating models) be meaningfully contrasted to drive the development of theory. Adopting the framework above, which is based in established frameworks for statistical decision theory (Savage, 1972) and information design in economics (Bergemann and Morris, 2019), also stands to improve researchers’ awareness of the limitations on what be concluded from study results when a decision problem is under-communicated.
The variation induced when a decision problem is not communicated is consequential to interpretation in that it cannot be separated from variation induced by true differences in how people do a task under different conditions. Identifying the key sources of loss—much less separating them given experiment results—is a hard problem made impossible when the researcher has not clearly identified the decision problem to the study participant or the reader of a resulting research paper. Consequently, researchers wishing to design a normative decision experiment should communicate all necessary components of the decision problem both to study participants and in publications. Specifically, we recommend that researchers:
@@
Provide prior and/or sufficient information about the data-generating model to calculate posterior. Sufficient information about the data-generating model must be provided to arrive at the Bayesian posterior distribution. Whenever the signal on its own does not induce a posterior, the prior should be endowed (Equation 4), and sufficient information about should be provided to enable estimating and . Using presentation formats that are more likely to be automatically processed (e.g., conveying distribution information via frequency animations (Goldstein and Rothschild, 2014; hullman2015hops)) or attempting to elicit participants’ priors directly (e.g., (Kim et al., 2019, 2020)) may help researchers minimize prior loss.
Explicitly motivate scoring rules for incentivization and evaluation. Beyond defining the components of the target decision problem, a next step is for researchers to motivate the problem definition they use. Recent work describes the use of a rational agent framework similar to the one presented here to analyze whether a decision study is likely to provide agents with sufficient incentive to consult the signals (Wu et al., 2023). Theory around scoring rules (Li et al., 2022) and what statistical quantities are elicitable (Lambert, 2011) stand to offer prescriptive advice if incorporated in human decision research.
Compare to best attainable performance. We observed that even when researchers invest effort to design scoring rules to incentivize participants, they do not often compare the observed behavioral performance to the best attainable performance under the rule. Researchers should default to analyzing results in the space of the scoring rule used for incentives, and provide a rationale whenever they deviate. Unbiased loss estimation is a rich area for future work.
@@
Conduct design analysis. Researchers doing normative decision research can benefit from conducting analyses prior to running the study which simulate the idealized actions of an agent to vet the experiment design (see, e.g., Wu et al. (2023)’s rational agent-based pre-study analyses). Simulation-based design analysis forces the researcher to explicitly define the parameters of their decision problem, which can elucidate what information needs to be communicated to a participant in order to enable an optimal response. It also enables the researcher to test variations on their signaling strategy and scoring rule, which can help the researcher realize under what conditions (1) the experiment can have acceptable statistical power and (2) participants can be paid fairly. It helps the researcher choose conditions (e.g., AI explanations) that provide informative signals for the decision task, narrowing down what kinds of research questions can be answered using a particular decision problem.
6. Limitations & Future Work
@@
Several challenges have been raised against expected utility theory as put forth by Savage (Savage, 1972); we recommend (Steele and Stefánsson, 2015) for an overview. Another critique of a normative approach dubbed “value collapse” (Nguyen, 2023) is that presenting a clear set of values as a target for behavior limits variance in ways that can be harmful. Our argument is that we should not mistake the variation that emerges from in an under-communicated decision problem as representative of behavior under a particular set of values or a particular real-world decision task. Studies aimed at understanding the range of values at play in a naturalistic decision scenario are better characterized as descriptive research. We note that while our framework is intended for normative decision research, the definitions below may also be useful if one is doing descriptive or exploratory decision research, by making clear which aspects of the decision problem are underspecified and therefore presumably the object of study. For example, a researcher might aim to characterize challenges in model-assisted decisions in a particular domain (Zytek et al., 2021), or conduct an exploratory study to elicit information about the utility function (i.e., scoring rule) that operates in a setting then communicate it back to new participants as part of the decision problem in an evaluative experiment (e.g., (Fernandes et al., 2018)).
@@
Some have voiced pessimistic views on the value of monetary incentives for experiment participants (e.g., (Read, 2005)). These points are somewhat orthogonal to the crux of our argument, which is philosophical and concerned with what is knowable more so than the pragmatics of how much behavior changes when a problem is well-defined.
@@ Acknowledgements.
We thank Yifan Wu and Ziyang Guo for discussions on these topics, and Dan Weld for feedback on a draft.References
@@
- Gneiting (2011) Tilmann Gneiting. 2011. Making and evaluating point forecasts. J. Amer. Statist. Assoc. 106, 494 (2011), 746–762.
@@
- Gonzalez et al. (2020) Ana Valeria Gonzalez, Gagan Bansal, Angela Fan, Robin Jia, Yashar Mehdad, and Srinivasan Iyer. 2020. Human evaluation of spoken vs. visual explanations for open-domain qa. arXiv preprint arXiv:2012.15075 (2020).
@@
- Martínez-Marquina et al. (2019) Alejandro Martínez-Marquina, Muriel Niederle, and Emanuel Vespa. 2019. Failures in contingent reasoning: The role of uncertainty. American Economic Review 109, 10 (2019), 3437–3474.
@@
- Wang and Yin (2021) Xinru Wang and Ming Yin. 2021. Are explanations helpful? a comparative study of the effects of explanations in ai-assisted decision-making. In 26th international conference on intelligent user interfaces. 318–328.
@@
- Yang et al. (2020) Fumeng Yang, Zhuanyi Huang, Jean Scholtz, and Dustin L Arendt. 2020. How do visual explanations foster end users’ appropriate trust in machine learning?. In Proceedings of the 25th international conference on intelligent user interfaces. 189–201.

### v5

@@ 6. Recommendations for Practice
The lack of consensus around how to define a decision problem and the minimum required components for normative analysis leads to research landscape in which researchers can easily be tempted to overinterpret the results of their studies. Readers may not recognize the overinterpretations, leading to a research literature based on invalid claims about human biases. Adopting a more rigorous standard in the form of the proposed framework, based in established theoretical foundations for decision making from informative signals (Savage, 1972; Bergemann and Morris, 2019), stands to improve empirical decision research in several ways. First, it improves the validity of interpretations of individual studies. The variation induced when a decision problem is not clearly reported nor communicated to participants is consequential to interpretation in that it cannot be separated from variation induced by true differences in how people do a task under different conditions. Identifying the key sources of loss—much less separating them given experiment results—is a hard problem made impossible. Adopting a standard where the normative decision arises from a well-defined process stands to improve researchers’ ability to point to possible sources of performance loss, and raise their awareness of the limitations on what be concluded from study results when a decision problem is under-communicated. Second, it paves the way for more rigorous theory development. Only when researchers define the task they study as a complete decision problem can the results of experiments that study slightly different problems (e.g., use different data-generating models) be meaningfully contrasted to drive the development of theory.
Consequently, researchers wishing to design a normative decision experiment should communicate all necessary components of the decision problem both to study participants and in publications. Specifically, we recommend that researchers:
@@
Provide prior and/or sufficient information about the data-generating model to calculate posterior. Explicit specification of priors is rare in decision studies used to evaluate interfaces in practice. As discussed above, sufficient information about the data-generating model must be provided to arrive at the Bayesian posterior distribution. Whenever the signal on its own does not induce a posterior, the prior should be endowed with sufficient information about should be provided to enable estimating and . Exemplar-based training procedures such as have been used to teach humans when to defer to an AI model can be employed for this purpose . More generally, using presentation formats that are more likely to be automatically processed (e.g., conveying distribution information via frequency animations (Goldstein and Rothschild, 2014; Hullman et al., 2015)) or validating participants’ priors through elicitation (e.g., (Kim et al., 2019, 2020)) can help researchers understand and minimize prior loss.
Explicitly motivate scoring rules for incentivization and evaluation. Beyond defining the components of the target decision problem, a next step is for researchers to motivate the problem definition they use. Recent work describes the use of a rational agent framework similar to the one presented here to analyze whether a decision study is likely to provide agents with a sufficient incentive to motivate them to consult the signals (Wu et al., 2023). Theory around scoring rules (Li et al., 2022) and what statistical quantities are elicitable (Lambert, 2011) stand to offer prescriptive advice if incorporated in human decision research.
Compare to best attainable performance. We observed that even when researchers invest effort to design scoring rules to incentivize participants, they do not often compare the observed behavioral performance to the best attainable performance under the rule. Researchers should default to analyzing results in the space of the scoring rule used for incentives, and provide a rationale whenever they deviate. How to estimate sources of loss in an unbiased way is a rich area for future work.
@@
Conduct design analysis. Researchers doing normative decision research can benefit from conducting analyses prior to running the study which simulate the idealized actions of an agent to vet the experiment design (see, e.g., Wu et al. (2023)’s pre-study analyses for visualization studies and Guo et al. (2024)’s analysis of the potential for complementarity in AI-assisted decisions). Simulation-based design analysis forces the researcher to explicitly define the parameters of their decision problem, which can elucidate what information needs to be communicated to a participant in order to enable an optimal response. It also enables the researcher to test variations on their data-generating model and scoring rule, which can help the researcher realize under what conditions (1) the experiment can have acceptable statistical power and (2) participants can be paid fairly. Design analysis helps the researcher choose conditions (e.g., AI explanations) that provide informative signals for the decision task, narrowing down what kinds of research questions can be answered using a particular decision problem. For example, at a minimum, to be informative of how effectively people use signals, a study design must have sufficient statistical power to separate the rational agent performance with and without the signal.
6.1. Discussion
By arguing that researchers adopt and sufficiently communicate to participants a widely-applicable definition of a decision problem, our goal is to improve the integrity of claims made from human subjects experiments. We now address how the proposed normative decision framework sheds light on fundamental uncertainty in generalizing study results to real-world contexts. We discuss several common misunderstandings of a rational agent approach to defining normative decisions. We conclude with limitations and directions for future work.
6.2. Goals and Incentives in the Lab versus the World
Imagine that the researchers who designed the voting task intentionally did not include the vote in the scoring rule, and were intentionally ambiguous about the role of the vote in the data-generating model. Because the utility of voting in a real-world election is typically known only implicitly at an individual level and likely to vary between people, the researchers might think that their results would therefore be more representative of real world voting behaviors.
Consequently, when people are in an experiment, we should not expect the analogous real world goals and beliefs for that kind of task to apply. For example, people might take actions in the real world for intrinsic value—e.g., voting in order to feel like a good citizen, or paying close attention to features of a flight they are deciding whether to book because the decision will impact whether or not they arrive in time for an important event. It is difficult to motivate people to take actions based on intrinsic value in an experiment, unless the experiment is designed to elicit and describe those intrinsic values, in which case the normative framework here no longer applies. Hence, the idea that as experimenters we can skip fully describing the decision problem to our participants while still retaining the ability to identify their behavior as faulty or biased arises from a failure to recognize our fundamental uncertainty about how the experimental context relates to the real world.
This requirement of communicating to study participants sufficient information for a rational agent to optimize may seem unnecessarily strict if one assumes that participants’ domain knowledge alone should enable them to evaluate the information about the state captured in signals. For example, in the flight booking example, a researcher may be tempted to assume that even without being given any information about the probability that the model prediction is correct, a participant will be able to use their prior knowledge about features of delayed flights to infer the probability that the model prediction is correct. The problem is that once a researcher makes this assumption, they are no longer evaluating participants on only the information they provided. They are also implicitly evaluating them on how well their prior beliefs about the state and the validity of the model’s learned information align with the researcher’s choice of data-generating model. There are various implications of this misalignment that we may find problematic upon reflection. For example, under such conditions, participants must bring identical beliefs to the researchers’ in order to have a chance of scoring well.
Fortunately, this does not mean that human decisions can only be evaluated by assuming that participants bring no additional information. Instead, the data-generating model and decision task should account for the presence of human prior information. For example, for AI-assisted decision tasks like the flight booking example (2.3), a human prediction of the delay status of a flight, governed by a distinct generating process from the AI prediction, can be included in the data-generating process. The decision problem for which the normative decision is sought can then be cast as a decision between the human and AI prediction. If independent human predictions are not available, they can be elicited from the subjects before providing the signal. Recent work by Guo et al. (Guo et al., 2024) demonstrates this approach.
6.3. Common Misunderstandings
Our argument is that the interpretation of results is restricted when a well-defined decision problem is not conveyed to participants, not that researchers cannot conduct comparisons at all without a well-defined decisions. Researchers can conduct and report on studies that do not conform to this approach, but they cannot conclude that people made bad decisions if they were not given sufficient information to know what a good decision is. For example, imagine a study comparing the decision performance of users with and without the use of an AI, where no prior is given to participants, and no scoring rule is given. Imagine that performance in the study is measured in terms of accuracy, and average accuracy is higher among participants who had access to the AI. Given an appropriately designed randomized experiment, the researcher could conclude that there is a causal effect of the AI on decisions. However, they could not claim that people are making worse or biased decisions without the AI, because they are scoring their decisions using information they never provided to participants. When we say that one condition did better, we should acknowledge that we did not give them a chance to fully understand and optimize their behavior for the problem that we analyzed.
Finally, a common misinterpretation of the use of a rational agent model is that we are implying that human behavior is expected to be perfectly rational, or approximately rational. This is not the case. While a rational agent approach formally justifies what it means to best use information to inform the selection of an action using statistical decision theory, participants in our experiment may use different strategies. The point is that without a definition of what it means to best respond to the information we have provided, we are hard-pressed to evaluate participant behavior. Even if we do not expect this procedure to characterize what human agents do, it provides a basis for interpreting the results in terms of deviations from a well-understood standard.
6.4. Limitations & Future Work
Several challenges have been raised against expected utility theory as put forth by Savage (Savage, 1972); we recommend (Steele and Stefánsson, 2015) for an overview. Another critique of a normative approach dubbed “value collapse” (Nguyen, 2023) is that presenting a clear set of values as a target for behavior limits variance in ways that can be harmful. Our argument is that we should not mistake the variation that emerges from in an under-communicated decision problem as representative of behavior under a particular set of values or a particular real-world decision task. If a researcher wishes to evaluate decisions but contests Bayesian decision theory as a normative standard, they should be prepared to motivate whatever standard they adopt instead with equivalent rigor (e.g., derivation from interpretable assumptions combined with mathematical laws). We note that alternative models have been well-explored, but rarely achieve provable grounding relative to mathematical laws as statistical decision theory does.
@@ 6.4. Limitations & Future Work
Sometimes researchers are explicitly interested in understanding the range of values at play in a naturalistic decision scenario rather than providing a well-defined decision problem. In this case, their research is better characterized as descriptive, and they should refrain from claims that hinge on evaluating participant decisions. While our framework is intended for normative decision research, the definitions may also be useful if one is doing such descriptive or exploratory decision research, by making clear which aspects of the decision problem are underspecified and therefore presumably the object of study. For example, a researcher might aim to characterize challenges in model-assisted decisions in a particular domain (Zytek et al., 2021), or conduct an exploratory study to elicit information about the utility function (i.e., scoring rule) that operates in a setting and then communicate it back to new participants as part of the decision problem in an evaluative experiment (e.g., (Fernandes et al., 2018)).
@@
Some have voiced pessimistic views on the value of monetary incentives for experiment participants (e.g., (Read, 2005)). As Wu et al. (2023) show, this view is to be expected under a rational agent approach whenever the value of information to the decision problem (i.e., the difference in the rational agent’s expected score given only the prior versus with access to the signals (Wu et al., 2023)) is small. More broadly, the insensitivity of human participants to monetary incentives are orthogonal to the crux of our argument, which is philosophical and concerned with what is knowable more so than the pragmatics of how much behavior changes when a problem is well-defined.
@@ Acknowledgements.
We thank Yifan Wu and Ziyang Guo for discussions on these topics, and Dan Weld and Matt Groh for feedback on a draft.References
@@
- Alur et al. (2024) Rohan Alur, Manish Raghavan, and Devavrat Shah. 2024. Distinguishing the Indistinguishable: Human Expertise in Algorithmic Prediction. arXiv preprint arXiv:2402.00793 (2024).
@@
- Corvelo Benz and Rodriguez (2024) Nina Corvelo Benz and Manuel Rodriguez. 2024. Human-aligned calibration for ai-assisted decision making. In Advances in Neural Information Processing Systems, Vol. 36.
@@
- Guo et al. (2024) Ziyang Guo, Yifan Wu, Jason D Hartline, and Jessica Hullman. 2024. A Decision Theoretic Framework for Measuring AI Reliance. In The 2024 ACM Conference on Fairness, Accountability, and Transparency. 221–236.
- Hofman et al. (2020) Jake M Hofman, Daniel G Goldstein, and Jessica Hullman. 2020. How visualizing inferential uncertainty can mislead readers about treatment effects in scientific results. In Proceedings of the 2020 chi conference on human factors in computing systems. 1–12.
- Hullman et al. (2015) Jessica Hullman, Paul Resnick, and Eytan Adar. 2015. Hypothetical outcome plots outperform error bars and violin plots for inferences about reliability of variable ordering. PloS one 10, 11 (2015), e0142444.